OpenStudy (anonymous):

Solve the following equation. 2/3x +6= 1/2x + 1/4x 2 6 14 2/5 72

OpenStudy (john_es):

Find the least common factor, 12, and rewrite, \[\frac{8}{12}x+\frac{72}{12}=\frac{6}{12}x+\frac{3}{12}x\] \[8x+72=6x+3x\] And solve for x. The solution should be x=72.

OpenStudy (anonymous):

thanks:)

OpenStudy (john_es):

;)